OpenStudy (anonymous):

The polynomial below is a perfect square trinomial of the form A2 - 2AB + B2

OpenStudy (anonymous):

9x to the second power - 30x + 25

OpenStudy (anonymous):

I'm stuck how do I figure this out?:/

OpenStudy (campbell_st):

ok... the simple way to do this is look at the 1st and last terms 9x^2 = (3x)^2 and 25 = 5^2 so A = 3x and B = - 5 the reason B is negative is that the middle term is negative... so is -30x = -2AB you need to decide... \[-30x = 2 \times 3x \times -5\]